OpenStudy (anonymous):

I am needing to get my last expression on a perfect square v^2-18+?

OpenStudy (domebotnos):

Any choices? Because +81 would work.

OpenStudy (domebotnos):

And don't forget to put your v on the middle term. ;) -18v

OpenStudy (solomonzelman):

\(\Large\color{black}{ \bf v^2-18v=0 }\) \(\Large\color{black}{ \bf v^2-18v+81=0 }\) \(\Large\color{black}{ \bf (v-9)^2=0 }\) \(\Large\color{black}{ \bf v=~? }\)